solve using any method (substitution, addition/subtraction, or multiplication). show work
x=-2y
3x+y=10

Answer by robertb(5830)

Substitute the top equation x = -2y into the bottom equation: 3(-2y) + y = 10.
-6y + y = 10,
-5y = 10,
y = -2. and x = -2*-2 = 4.